# Bridged Clustering for Representation Learning: Semi-Supervised Sparse Bridging

### v1

@@ 4 Bridged Clustering Algorithm
On the $k$ paired points, we match input-space clusters to output-space clusters by majority vote. For every input cluster $a$, mapping $\hat{A}(a)\;:=\;\operatorname*{arg\,max}_{b\in[C]}\sum_{(x^{\prime},y^{\prime})\in\mathcal{S}}\bm{1}\!\{\hat{c}_{X}(x^{\prime})=a,\;\hat{c}_{Y}(y^{\prime})=b\}.$ For a random latent label $T$, we call $\varepsilon_{B}:=\Pr\bigl[\hat{A}(T)\neq T\bigr]$ the mis-bridging rate. Algorithm 1 predicts a fresh output by returning the empirical centroid of the output cluster indicated by $\hat{A}\circ\hat{c}_{X}$. Let $\widehat{\mu}_{b}^{Y}$ be the empirical mean of $\mathcal{Y}$-points assigned to cluster $b$, then $\hat{Y}(x):=\widehat{\mu}_{\hat{A}(\hat{c}_{X}(x))}^{\,Y}.$
5 Algorithmic Analysis
This section provides a theoretical understanding of Bridged Clustering, identifying key factors governing its algorithmic effectiveness and efficiency.
5.1 Inference Effectiveness
For a random sample $(X,Y)\;\big|\;T$, the Bridged Clustering Algorithm is tasked with retrieving an estimation of $Y$ based on $X$. To analyze the effectiveness of inference, we atomize the process:
$X\xrightarrow[\varepsilon_{X}]{\hat{c}_{X},(\sigma_{X})}\hat{T}\xrightarrow[\varepsilon_{B}]{\hat{A}}\hat{\mu}_{\hat{T}}^{Y}=\hat{Y}\xrightarrow[\varepsilon_{Y},D_{Y}]{\hat{c}_{Y},(\sigma_{Y})}Y$
Our model first implicitly predicts a latent label $\hat{T}$ for the sample input $X$ by using $\hat{c}_{X}$ to assign it to the closest input cluster. Then, using the learned cluster-to-cluster mapping $\hat{A}$, our model reaches the corresponding output cluster centroid $\hat{\mu}_{\hat{T}}^{Y}$. From there, the distance of our estimated output $\hat{Y}$ from the actual output $Y$ is mainly determined by two factors: whether $Y$ is assigned by $\hat{c}_{Y}$ to the correct output cluster $T$ (bounded by $\varepsilon_{Y}$), and—if the assignment is correct—the in-cluster distance between the $Y$ and its cluster centroid; we define this distance as $D_{Y}$.
Throughout the process, $\hat{A},\hat{c}_{X},\hat{c}_{Y}$ are the key functions solidified during model training, learned from the supervised data, unsupervised inputs, and unsupervised outputs, respectively. The success of each step during inference is quantified through $\varepsilon_{X},\varepsilon_{Y},\varepsilon_{B},D_{Y}$. Correspondingly, as our model becomes robust in input-assignment, output assignment, and bridging accuracy, the $\varepsilon_{X},\varepsilon_{Y},\varepsilon_{B}$ are actively controlled so that the model makes more accurate inferences.
5.2 Risk Control
We quantify how representation quality and scarce supervision govern the three error sources in Bridged Clustering: input mis-clustering $(\varepsilon_{X})$, output mis-clustering $(\varepsilon_{Y})$, and mis-bridging $(\varepsilon_{B})$. The high-level takeaways are that (1) stronger embeddings that enlarge inter-clustering separation directly reduce $\varepsilon_{X}$ and $\varepsilon_{Y}$, and (2) once clusters are reliable, even a small handful of paired examples suffice to learn the sparse cluster bridge with exponentially small error.
Mis-clustering under sub-Gaussian mixtures.
Suppose the input and output distributions are $\sigma^{2}$–sub-Gaussian with means $\mu_{t}^{X},\mu_{t}^{Y}$. Define the minimum separation $\Delta_{X}=\min_{t\neq s\in[C]}\|\mu_{t}^{X}-\mu_{s}^{X}\|$ and $\Delta_{Y}=\min_{t\neq s\in[C]}\|\mu_{t}^{Y}-\mu_{s}^{Y}\|$. Classical results on Lloyd’s algorithm show that, under good initialization, $C$-means clustering on $n_{X}$ unlabeled inputs satisfies
with probability at least $1-\nu-\frac{4}{n_{X}}-2\exp(-\Delta_{X}/\sigma)$, where $\nu$ is the initializer failure rate (Lu & Zhou, 2016). An identical bound holds for $\varepsilon_{Y}.$
Thus, any improvement in embeddings that increases $\Delta_{X}$ or $\Delta_{Y}$ sharpens the bound without changing the algorithm. Consistent with this theory, we observe a strong negative correlation between cluster quality (AMI) and predictive loss (Appendix E).
Efficient bridging from few pairs.
Once the mis-clustering rates $\varepsilon_{X},\varepsilon_{Y}$ are controlled, learning the bridge becomes simple. For a supervised pair $(X,Y)|T\in\mathcal{S}$, consider the events $\{\sigma_{X}(\hat{c}_{X}(X))=T\}$ and $\{\sigma_{Y}(\hat{c}_{Y}(Y))=T\}$, which correspond to correct cluster assignments in the input and output spaces, controlled via $\varepsilon_{X},\varepsilon_{Y}$. If both events hold with sufficiently high probability $(>50\%)$, then a majority vote over just a few labeled pairs will suffice to recover the correct cluster-to-cluster assignment, resulting in a low mis-bridging error $\varepsilon_{B}$. Empirically, we find that even one or two examples per cluster are enough for reliable bridging, explaining why Bridged Clustering remains effective with less than $1\%$ paired supevision (Sec. 6).
5.3 Runtime and Memory Complexity
$\text{training}=f(n_{X},C,d)+f(n_{Y},C,d^{\prime})+O(k),\qquad\text{per-query inference}=O(Cd+d^{\prime}),$
Method Training time Memory
EOT $O(S\,n_{X}n_{Y})$ $O(n_{X}n_{Y})$
GW $O(T_{\text{GW}}(n_{X}^{2}n_{Y}+n_{X}n_{Y}^{2}))$ $O(n_{X}^{2}+n_{Y}^{2}+n_{X}n_{Y})$
@@ 6 Experiments
We now turn to empirically test the predictive accuracy of Bridged Clustering. We use fully paired datasets with known input-output correspondences ($(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}$). These datasets allow us to simulate the disjoint setting by hiding most of the pairings and exposing only a small, randomly selected subset during training. This controlled setup enables us to measure performance as a function of the number of labeled pairs, while retaining access to ground-truth labels for evaluation.
@@ 6.1 Data
We apply Bridged Clustering on four multimodal datasets that connect distinct feature spaces: the BIOSCAN-5M dataset that pairs specimen images with DNA barcode sequences of insect samples (Gharaee et al., 2024), the WIT dataset with Wikipedia Image-Text pairings (Srinivasan et al., 2021), as well as the Flickr30k dataset (Young et al., 2014) and the COCO dataset (Lin et al., 2014) that connect everyday images with human-generated descriptions. We randomly sample a number of data groups to include in each run of our experiments, e.g., selecting 5 insect species from Bioscan, or 7 object categories from COCO, etc. We detail the data, processing, and mapping in Appendix A.
@@ 6.2 Experimental Overview
We evaluate Bridged Clustering’s ability to predict the output features $\mathcal{Y}$ (e.g., DNA sequences or text captions) from the input features $\mathcal{X}$ (e.g., images). For each dataset, we simulate a disjoint data setting by designating a small subset of datapoints as the supervised set, which contains paired input-output examples. The remaining datapoints are divided into two unpaired sets: one with input-only features and the other with output-only features, mimicking independent data sources. During training, the model leverages 3 sources of information: (1) unpaired input features ($\mathcal{X}$), (2) unpaired output features ($\mathcal{Y}$), and (3) a small set of paired examples $(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}$. For evaluation, we assess the model’s ability to recover the missing output features $\mathcal{Y}$ of the input-only datapoints $\mathcal{X}$.
One of the advantages of Bridged Clustering is its bidirectionality. Thus, in addition to predicting $\mathcal{Y}$ from $\mathcal{X}$, we also test the inverse by predicting input features $\mathcal{X}$ from output features $\mathcal{Y}$.
We compute the Mean Squared Error (MSE) between the predicted and the ground-truth output features. In image captioning tasks, outputs are text embeddings; thus, we map each predicted embedding to its nearest neighbor in the unpaired output set. We then compute the MSE between this retrieved embedding and the ground-truth caption embedding (see details in Appendix A).
@@ 6.3 Baselines
We compare Bridged Clustering against a comprehensive suite of baselines spanning supervised learning, classical transductive methods, graph-based models, consistency-regularized deep learners, pairwise regression approaches, and unmatched regression techniques. Together, they span the full landscape of SSL and related paradigms. All except KNN leverage unlabeled inputs and paired data. Bridged Clustering, KMM, EM, EOT, GW incorporate unlabeled outputs as well.
@@
  • Graph-structured: Graph Convolutional Network (GCN).
@@ 6.4 Results
We evaluate Bridged Clustering on four datasets, testing its ability to map $x\in\mathcal{X}$ to $y\in\mathcal{Y}$ and vice versa. Since Bridged Clustering assumes data can be grouped discretely, each experiment samples a subset of groups (e.g., species in Bioscan). We test the robustness of Bridged Clustering by sampling different numbers of data groups $\in\{3,4,5,6,7\}$ and varying the numbers of supervised samples per cluster $\in\{1,2,3,4\}$ [^3] Corresponds to $\{0.5\%,1\%,1.5\%,2\%\}$ supervision for Bioscan and COCO (cluster size=200), and $\{4\%,8\%,12\%,16\%\}$ for WIT and Flickr30k (cluster size=25)., yielding $4\times 5=20$ different settings per experiment.
Each setting is run 30 times with different seeds. Figure 1 shows MSE distributions, while Table 1 reports win-rates (lowest MSE across 600 runs per dataset). Additional results in Figures A.2-A.5. Overall, Bridged Clustering outperforms all baselines across all datasets except WIT. We note the robustness of the results across transductive and inductive setups.
Figure 1: MSE distribution of different models in the inductive setting. The four distribution plots of the same color represent the settings with 1, 2, 3, and 4 supervised samples per cluster.
@@
Exp. BC KNN FM LapR. TSVR TNNR UCVME GCN KMM EM EOT GW BIOSCAN 0.67 0.00 0.00 0.00 0.19 0.00 0.00 0.00 0.13 0.01 0.01 0.00 BIOSCAN (rev.) 0.63 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.01 0.01 0.36 0.00 COCO 0.71 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.03 0.26 0.00 COCO (rev.) 0.89 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.04 0.07 0.00 Flickr30k 0.56 0.00 0.00 0.00 0.00 0.04 0.00 0.00 0.01 0.02 0.36 0.01 Flickr30k (rev.) 0.71 0.00 0.00 0.00 0.13 0.00 0.00 0.00 0.03 0.00 0.10 0.04 WIT 0.12 0.00 0.00 0.00 0.00 0.32 0.00 0.00 0.28 0.03 0.16 0.10 WIT (rev.) 0.09 0.00 0.00 0.00 0.04 0.30 0.00 0.00 0.27 0.00 0.05 0.25
@@
Exp. BC KNN FM LapR. TSVR TNNR UCVME GCN KMM EM EOT GW BIOSCAN 0.67 0.00 0.00 0.00 0.18 0.00 0.00 0.00 0.11 0.01 0.03 0.00 BIOSCAN (rev.) 0.61 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.01 0.01 0.37 0.00 COCO 0.70 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.03 0.28 0.00 COCO (rev.) 0.84 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.05 0.11 0.00 Flickr30k 0.45 0.01 0.00 0.00 0.00 0.04 0.00 0.00 0.01 0.03 0.46 0.01 Flickr30k (rev.) 0.60 0.01 0.00 0.00 0.00 0.01 0.00 0.00 0.03 0.01 0.32 0.03 WIT 0.12 0.00 0.00 0.00 0.00 0.32 0.00 0.00 0.28 0.03 0.16 0.10 WIT (rev.) 0.09 0.00 0.00 0.00 0.04 0.30 0.00 0.00 0.27 0.00 0.05 0.25
Table 1: Win-rates across transductive (a) and inductive (b) experiments. Bold indicates the best model per dataset. Win-rates are defined as the model with the lowest MSE across 600 runs in every dataset (30 randomized trials per setting). For example, if Bridged Clustering achieves the lowest MSE among all models in 300 out of the 600 trials for some setting, its score is 0.50.
7 Discussion
Bridged Clustering outperforms baseline performance in three out of four datasets (we will discuss WIT’s performance shortly). In these datasets, Wilcoxon signed-rank tests confirmed these findings are statistically significant (largest p-value $6\mathrm{e}{-4}$).
@@ 7 Discussion
Across both directions ($\mathcal{X}\rightarrow\mathcal{Y}$, $\mathcal{Y}\rightarrow\mathcal{X}$) in Bioscan, Flickr30k, and COCO, Bridged Clustering achieves consistently low MSE, outperforming all baselines over varied cluster sizes and supervision levels. On WIT, it falls short of the best model but remains highly competitive, notably surpassing all SSL baselines by leveraging output-only data $\mathcal{Y}$. On average, Bridged Clustering also outperforms KMM, EM, EOT, and GW, despite these methods using full training data, underscoring the strength of our simpler approach. While transport-based models are strong baselines, Bridged Clustering offers greater interpretability and efficiency.
@@
Bridged Clustering’s performance depends on recovering latent structure via independent clustering (cluster quality) and aligning clusters with limited supervision (bridging accuracy). Theoretically, inference quality is bounded by mis-clustering rates $\varepsilon_{X},\varepsilon_{Y}$ and bridging accuracy $\varepsilon_{B}$, which together limit model risk. Empirically, cluster quality (adjusted mutual information) negatively correlates with model loss (Figure A.8). WIT’s low mutual information suggests that weaker clustering explains its lower performance (Appendix E).
@@
Bridged Clustering performs well in low-supervision settings, as it requires as few as one labeled example per cluster to associate input and output clusters accurately. While increasing the number of supervised points can improve alignment accuracy, it has a limited effect on prediction quality. In other words, since the model always returns the output cluster centroid, further supervision does not refine individual predictions. In contrast, other models (excluding EOT) may benefit more from additional supervision.
@@
Bridged Clustering’s relative weakness in high-cluster regimes stems from its reliance on bijection input-output mapping. As cluster count increases, the space of possible mappings grows exponentially, making accurate alignment harder (See Figure 11 in Appendix E). Baseline methods without such explicit mappings prove more robust to increases in cluster count, highlighting Bridged Clustering’s tradeoff: it excels in low-supervision, low-cluster settings, but becomes more sensitive as structural complexity increases.
Data-efficient learning:
Bridged Clustering demonstrates that modular architectures leveraging unsupervised structure and lightweight alignment can match or exceed fully supervised models. By separating representation learning from supervision, it uncovers latent structure in input and output spaces and aligns them with minimal supervision. This approach is especially effective with limited labeled data, high domain variability, or distribution shifts, where monolithic models often overfit. Our results indicate that concept-level alignment between independently clustered spaces provides a strong inductive bias, enabling generalization without dense supervision.
Conceptual appeal:
By decoupling structure discovery (unsupervised clustering) from cross-space association (a small, sparse bridge), Bridged Clustering offers a simple blueprint for learning with abundant unpaired observations. This model-agnostic algorithm can be extended (hierarchical clusters, soft bridges, learned metrics) while preserving interpretability and efficiency.

### v3

@@ 4 Bridged Clustering Algorithm
On the $k$ paired points, we match input-space clusters to output-space clusters by majority vote. For every input cluster $a\in[C]$, mapping
$\hat{A}(a)\;:=\;\operatorname*{arg\,max}_{b\in[C]}\sum_{(x^{\prime},y^{\prime})\in\mathcal{S}}\mathbf{1}\!\{\hat{c}_{X}(x^{\prime})=a,\;\hat{c}_{Y}(y^{\prime})=b\}.$
For a random latent label $T$, we call
$\varepsilon_{B}:=\Pr\!\left[\sigma_{Y}\!\left(\hat{A}(\sigma_{X}^{-1}(T))\right)\neq T\right].$
the mis-bridging rate. Algorithm 1 predicts a fresh output by returning the empirical centroid of the output cluster indicated by $\hat{A}\circ\hat{c}_{X}$. Let $\widehat{\mu}_{b}^{Y}$ be the empirical mean of $\mathcal{Y}$-points assigned to cluster $b$, so our final predictor is
$\hat{Y}(x):=\widehat{\mu}_{\hat{A}(\hat{c}_{X}(x))}^{\,Y}.$
@@ 5 Experiments
We now turn to empirically test the predictive accuracy of Bridged Clustering. We use fully paired datasets with known input-output correspondences ($(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}$) across various modalities. These datasets allow us to simulate independently collected $\mathcal{X}$-only and $\mathcal{Y}$-only repositories by hiding most of the pairings and exposing only a small, randomly selected subset during training. This controlled setup with ground-truth pairs enables us to measure performance as a function of the number of labeled pairs.
@@ 5.1 Data
We apply Bridged Clustering on four multimodal datasets that connect distinct feature spaces: the BIOSCAN-5M dataset that pairs specimen images with DNA barcode sequences of insect samples (Gharaee et al., 2024), the WIT dataset (Srinivasan et al., 2021) pairing images with their text captions on Wikipedia pages, as well as the Flickr30k dataset (Young et al., 2014) which connects images of people engaged in everyday activities with human-annotated captions and the COCO dataset (Lin et al., 2014) that connects common object images with text descriptions.
For experimental evaluation, we randomly sample a number of data groups, where each group corresponds to a coherent semantic category or label within the dataset (e.g., an insect species in BIOSCAN, or an object category in COCO). Each run includes a subset of these groups (e.g., 5 species of insects from BIOSCAN or 7 categories of objects from COCO) to assess Bridged Clustering’s ability to recover and align latent structure. We detail the data, processing, and mapping in Appendix A.
@@ 5.2 Experimental Overview
We evaluate Bridged Clustering’s ability to predict the output features $\mathcal{Y}$ (e.g., DNA sequences or text captions) from the input features $\mathcal{X}$ (e.g., images). Since Bridged Clustering is bidirectional, we also test the inverse by predicting input features $\mathcal{X}$ from output features $\mathcal{Y}$.
For each dataset, we simulate a disjoint data setting by designating a small subset of datapoints as the supervised set, which contains paired input-output examples. The remaining datapoints are divided into two unpaired sets: one with input-only features and the other with output-only features, mimicking independent data sources. We run experiments in two different settings: transductive, where the model has access to the unlabeled test inputs during training (with test outputs withheld), and inductive, where evaluation is performed on entirely unseen test data.
Metrics. We compute the Mean Squared Error (MSE) between the predicted and the ground-truth output features. For our BIOSCAN experiment, we directly measure the MSE between the algorithm-predicted embedding vector and the ground truth DNA embedding vector. In all three image captioning tasks (WIT/Flickr30k/COCO), we consider an additional metric that directly measures the accuracy of the predicted caption: BLEU score, besides the MSE. Since our outputs are text embeddings, we map each predicted embedding to its nearest neighbor in the available output set, which is a text embedding that corresponds to some actual text. We retrieve this text as our predicted caption. We then compute the BLEU score between this retrieved caption and the ground-truth caption. (see details in Appendix A).
@@ 5.3 Baselines
We compare Bridged Clustering against a comprehensive suite of baselines spanning supervised learning, classical transductive methods, graph-based models, consistency-regularized deep learners, pairwise regression approaches, and unmatched regression techniques (see more information in Section 2). Together, they span the full landscape of SSL and related paradigms, whereas all except KNN leverage unlabeled inputs and paired data, and Bridged Clustering, KMM, EM, EOT, GW incorporate unlabeled outputs as well. The following list categorizes the baselines we compare Bridged Clustering against:
@@
  • Graph-based: Graph Convolutional Network (GCN).
@@ 5.4 Results
Table 1: Win-rates across transductive (a) and inductive (b) experiments. Bold indicates the best model per dataset (each row is summed to 100%). Win-rates are defined as the model with the lowest MSE across 600 runs in every dataset (30 randomized trials per setting).
@@
Exp. BC KNN FM LapR. TSVR TNNR UCVME GCN KMM EM EOT GW BIOSCAN 67% 0% 0% 0% 19% 0% 0% 0% 13% 1% 1% 0% BIOSCAN (rev.) 63% 0% 0% 0% 0% 0% 0% 0% 1% 1% 36% 0% COCO 71% 0% 0% 0% 0% 0% 0% 0% 0% 3% 26% 0% COCO (rev.) 89% 0% 0% 0% 0% 0% 0% 0% 0% 4% 7% 0% Flickr30k 56% 0% 0% 0% 0% 4% 0% 0% 1% 2% 36% 1% Flickr30k (rev.) 71% 0% 0% 0% 13% 0% 0% 0% 3% 0% 10% 4% WIT 12% 0% 0% 0% 0% 32% 0% 0% 28% 3% 16% 10% WIT (rev.) 9% 0% 0% 0% 4% 30% 0% 0% 27% 0% 5% 25%
@@
Exp. BC KNN FM LapR. TSVR TNNR UCVME GCN KMM EM EOT GW BIOSCAN 67% 0% 0% 0% 18% 0% 0% 0% 11% 1% 3% 0% BIOSCAN (rev.) 61% 0% 0% 0% 0% 0% 0% 0% 1% 1% 37% 0% COCO 70% 0% 0% 0% 0% 0% 0% 0% 0% 3% 28% 0% COCO (rev.) 84% 0% 0% 0% 0% 0% 0% 0% 0% 5% 11% 0% Flickr30k 45% 1% 0% 0% 0% 4% 0% 0% 1% 3% 46% 1% Flickr30k (rev.) 60% 1% 0% 0% 0% 1% 0% 0% 3% 1% 32% 3% WIT 12% 0% 0% 0% 0% 32% 0% 0% 28% 3% 16% 10% WIT (rev.) 9% 0% 0% 0% 4% 30% 0% 0% 27% 0% 5% 25%
Each run of our experiments samples a subset of groups (e.g., species in BIOSCAN) from our datasets. We test the robustness of Bridged Clustering by sampling different numbers of data groups $\in\{3,4,5,6,7\}$ and varying the numbers of supervised samples per cluster $\in\{1,2,3,4\}$ [^2] Corresponds to $\{0.5\%,1\%,1.5\%,2\%\}$ supervision for BIOSCAN and COCO (cluster size=200), and $\{4\%,8\%,12\%,16\%\}$ for WIT and Flickr30k (cluster size=25)., yielding $4\times 5=20$ different settings per experiment.
Each setting is run 30 times with different seeds. Table 1 reports win-rates (lowest MSE across 600 runs per dataset), and Figure 2 shows MSE distributions. Additional results in Appendix Figures 4-9. Overall, Bridged Clustering outperforms all baselines across all datasets except WIT. We note the robustness of the results across transductive and inductive setups, and the fact that Bridged Clustering is always better in the transductive configuration.
Figure 2: MSE distribution of different models in the inductive setting. The four distribution plots of the same color represent the settings with 1, 2, 3, and 4 supervised samples per cluster. See more detailed MSE figures in the Appendix E.
5.5 Discussion
Bridged Clustering outperforms baseline performance in three out of four datasets. In these three datasets (we will discuss WIT’s performance shortly), Wilcoxon signed-rank tests confirmed these findings are statistically significant (largest p-value $6\mathrm{e}{-4}$).
@@ 5.5 Discussion
Across both directions ($\mathcal{X}\rightarrow\mathcal{Y}$, $\mathcal{Y}\rightarrow\mathcal{X}$) in BIOSCAN, Flickr30k, and COCO, Bridged Clustering achieves consistently low MSE, outperforming all baselines over varied cluster sizes and supervision levels. In WIT, it falls short of the best model, but remains highly competitive, notably surpassing all SSL baselines by leveraging output-only data $\mathcal{Y}$. On average, Bridged Clustering also outperforms KMM, EM, EOT, and GW, despite these methods using full training data, underscoring the strength of our simpler approach. Moreover, unlike transport-based methods, Bridged Clustering does not require solving costly optimal transport problems, leading to lower computational overhead, and produces easily interpretable cluster assignments aligned with conceptual groupings in the data (See Illustrations in Appendix J).
@@
Bridged Clustering’s performance depends on two factors: recovering latent structure via independent clustering (cluster quality) and aligning clusters with limited supervision (bridging accuracy). We observe that cluster quality is negatively correlated with model loss, measured using adjusted mutual information (AMI) (Figure A.10 in Appendix F). WIT’s low mutual information scores suggests that weaker clustering explains Bridged Clustering’s lower performance on this dataset. We also observe a positive correlation between bridging accuracy and cluster quality in 11 and 12, indicating that these two performance-relevant factors are correlated, a connection that’ll be rediscussed in Section 7.
@@
Bridged Clustering performs well in low-supervision settings, as it requires as few as one labeled example per cluster to associate input and output clusters accurately. While increasing the number of supervised points can improve alignment accuracy, it has a limited effect on prediction quality. We can see this in Appendix K.1, where Figures 23 - 28 show that lower supervision (1,2 sample) yields more competitive performance then higher supervision (3,4 sample) when compared to other methods, as measured by the task accuracy metric BLEU. In other words, since the model always returns the output cluster centroid, further supervision does not refine individual predictions. In contrast, other models may benefit more from additional supervision. To address the potential need for cluster-wise refinement, we designed an algorithmic variant in Section 6.2.
Transductive Versus Inductive:
The performance of Bridged Clustering across both transductive and inductive settings are consistently competitive, with slightly more wins in the transductive tests. Given that our clustering-based method naturally takes advantage of inference points to better learn the cluster geometry in the transductive settings, this marginal win is expected.
@@
Bridged Clustering’s relative weakness in high-cluster regimes stems from its reliance on bijection input-output mapping. As cluster count increases, the space of possible mappings grows exponentially, making accurate alignment harder (see Figure 13 in Appendix F). Baseline methods without such explicit mappings prove more robust to increases in cluster count, highlighting Bridged Clustering’s tradeoff: it excels in low-supervision, low-cluster settings, but becomes more sensitive as structural complexity increases.
6 Algorithmic Variants and Extended Applicability
To preserve the conceptual simplicity of our exposition, our main Algorithm 1 adopts general primitives: any clustering method for unsupervised learning and any mapping mechanism for sparse bridging.
To illustrate how Bridged Clustering could be adapted to more complex settings, we analyze two algorithmic variants, each addressing a range of data-centric challenges. In Sections 6.1 and 6.2, we describe these variants, and in Section 6.3, we describe the main takeaways of our empirical results using these variants under challenging data conditions.
6.1 Soft Bridging for Flexible Mapping
When applying our algorithm to datasets without clearly separated categories, clustering primitives may become prone to overlapping or miscounted clusters. Rigid one-to-one mapping, as in Section 4, becomes vulnerable to misattributed clusters under this setting.
Hence, we design a soft-bridging function as our new mapping primitive: using a small supervised set to build input-output cluster bridges, we construct probabilistic bridges between every possible input-output cluster pair, dynamically weighting each edge by the number of supervised points connecting the clusters. Formally, let $w_{i\to j}$ denote the fraction of supervised points whose inputs fall in input cluster $i$ and whose labels lie in output cluster $j$, with $\sum_{j}w_{i\to j}=1$. For an input $x$ assigned to cluster $i$, soft bridging predicts a weighted combination of all output cluster centroids:
$\hat{y}(x)\;=\;\sum_{j=1}^{C}w_{i\to j}\,\bar{y}_{j}.$
This method reduces the impact of enforcing a rigid one-to-one bridge when the clustering is imperfect. In practice, the clustering step may return the wrong number of clusters or be unable to distinguish overlapping clusters, which can lead to incorrect cluster correspondences. Soft bridging avoids committing to a single correspondence by distributing probability mass across output clusters, and our experiments (Section 6.3) show that it is robust on noisy data.
6.2 Cluster-wise Regression For Refined Prediction
Once a test point $x$ is assigned to an input cluster $i$ and bridged to an output cluster $j$, Algorithm 1 returns the output centroid $\widehat{\mu}_{j}^{Y}$, collapsing intra-cluster variation. To relax this rigidity, we propose cluster-wise regression, deployed as a third step after clustering and bridging. For each cluster pair, we use in-cluster supervised points to fit a linear map $f_{i\to j}(x)\in\mathbb{R}^{d^{\prime}}$ and predict
$\hat{y}(x)\;=\;(1-\alpha)\,\widehat{\mu}_{j}^{Y}\;+\;\alpha\,f_{i\to j}(x),$
with $\alpha\in[0,1]$ controlling the strength of the regression refinement. We instantiate two variants:
  • Supervised Refinement: linear regression on the supervised pairs in $(i,j)$;
  • Supervised+Centroid Refinement: linear regression on the supervised pairs plus the $C$ estimated input/output cluster centroids.
Cluster-wise regression is especially valuable as template for extending Bridged Clustering beyond very-low-supervision and conditionally independent settings. In regimes where the output $Y$ is conditionally linearly on the input $X$ given the clustered latent $T$, this offers a direct way to refine our cluster-based predictor, if provided sufficient labeled pairs to fit these cluster-wise regressors. Concretely, for a given $T=t$, if the relationship between $X$ and $Y$ is well approximated by a cluster-wise linear map $Y\approx W_{t}X+b_{t}$, we can reliably estimate that map with sufficient labeled pairs within each cluster.
6.3 Experiments with Challenging Data Conditions
In order to target overlapping clusters, we design a new experiment by sampling multi-labeled COCO classes that share at least one overlapping label with another sampled class, in contrast to our earlier use of single-labeled categories. Meanwhile, in order to target a misspecified number of clusters, we design another experiment in COCO by intentionally mis-specifying the number of clusters in our K-clustering primitives.
In the two experiments, both Soft-Bridging and Cluster-wise Regression (Supervised+Centroid Refinement) outperform Algorithm 1 under these challenging data settings, suggesting the adaptivity of Bridged Clustering as an effective template for diverse algorithms that jointly exploit unsupervised inputs and outputs. See Figure 14 through 17 in Appendix G and H.
7 Theoretical Discussion
This section provides the mathematical principals behind Bridged Clustering, identifying key factors governing its effectiveness. We also provide a runtime analysis in Appendix  D.
Representation quality and scarce supervision influence the final accuracy of Bridged Clustering through several key sources of error, including the input mis-clustering $(\varepsilon_{X})$, output mis-clustering $(\varepsilon_{Y})$, and mis-bridging $(\varepsilon_{B})$. The main takeaways of this section are that (1) stronger embeddings that enlarge inter-clustering separation directly reduce $\varepsilon_{X}$ and $\varepsilon_{Y}$, and (2) once clusters are reliable, even a small handful of paired examples suffice to learn the sparse cluster bridge with exponentially small error.
Mis-clustering under sub-Gaussian mixtures.
Suppose the input and output distributions are $\sigma^{2}$–sub-Gaussian with means $\mu_{t}^{X},\mu_{t}^{Y}$. Define the minimum separation
$\Delta_{X}=\min_{t\neq s\in[C]}\|\mu_{t}^{X}-\mu_{s}^{X}\|,\Delta_{Y}=\min_{t\neq s\in[C]}\|\mu_{t}^{Y}-\mu_{s}^{Y}\|.$
It is known that under Lloyd’s algorithm, under good initialization, $C$-means clustering on $n_{X}$ unlabeled inputs satisfies
with probability at least $1-\nu-\frac{4}{n_{X}}-2\exp(-\Delta_{X}/\sigma)$, where $\nu$ is the initializer failure rate (Lu and Zhou, 2016). An identical bound holds for $\varepsilon_{Y}.$ Thus, any improvement in embeddings that increases $\Delta_{X}$ or $\Delta_{Y}$ sharpens the bound without changing the algorithm. Consistent with this theory, we observe a strong negative correlation between cluster quality (AMI) and predictive loss (Appendix F).
Efficient bridging from few pairs.
Once the mis-clustering rates $\varepsilon_{X},\varepsilon_{Y}$ are controlled, learning the bridge becomes simple. As observed in our experiments, there’s a positive correlation between bridging accuracy and cluster quality. For a supervised pair $(X,Y)|T\in\mathcal{S}$, consider the events $\{\sigma_{X}(\hat{c}_{X}(X))=T\}$ and $\{\sigma_{Y}(\hat{c}_{Y}(Y))=T\}$, which correspond to correct cluster assignments in the input and output spaces, controlled via $\varepsilon_{X},\varepsilon_{Y}$. If both events hold with sufficiently high probability $(>50\%)$, then a majority vote over just a few labeled pairs will suffice to recover the correct cluster-to-cluster assignment, resulting in a low mis-bridging error $\varepsilon_{B}$. Empirically, we find that even one or two examples per cluster are enough for reliable bridging, explaining why Bridged Clustering remains effective with less than $1\%$ paired supervision (Section 5).